SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (9th Jan, 2007)

Attendees:

Operations

  • The old TSF fiber splitter has been removed and the fibers re-routed.
  • The old TSF crates and channels now need to be removed from EPICS. Walt has taken the crates out and Debbie has removed the 'old tsf' panel from the EPICS display, but they should be removed from the database and archiver as well.
  • The crate and boxes containing the old TSF need to be moved out of the electronics hut. Do they go to the TRG lab or can a home be found for them in IR2 somewhere?
  • The IFT timing jumper was moved - according to the email here just the bottom jumper needed to be moved. This made the IFT timing much better: see last plots of old and new. The IFT timing is still ~1 tick behind the other components: see first plot here. Now, the jumper labelled 'L2' was switched for Run 5 to account for 2 sectors being converted to LST (the plots given above have L2 switched) and it was thought perhaps the other jumpers could be switched in some combination to make a global timing change, see the e-logbook entries here. The current configuration is to run with L2 and the bottom jumper switched. It's still not clear what all these jumpers do, Maurizio will be asked. Also unclear whether it's OK to run like this.....
  • Trigger was paged on Friday, which turned out to be a problem with the new online release messing up database references, but only Debbie recieved the email. Need to change the 'on-call expert' in the contact db to be bbrtrg.

EMT

  • Dying ADB boards showed up in EMT fast mon plots (runs 69282 and 69283).
  • Looking at the config db - checking how to read global thresholds. See here, the conversion for ADC count to MeV is 0.4. The thresholds go: Y X E G M.

GLT

  • The work on the IFT turned up a bug that requires a GLT ROM reboot when a crate is powercycled. This was fixed somewhere along the way but has got lost. It needs to be fixed at high priority - it'll surely cause problems!

L3

  • The tracking group is yet to sign off on Arik's changes.

DQM

  • We need to get this put in the release before we get colliding beams - in ~2 weeks.

Config

  • Need to request MC samples with 1Zk included. Then, can compare simulated tight/emergency configurations with the emulated effect to check the emulator is working.
  • We need to confirm the charge is correct in the 1Zk rather urgently. Debbie will use fastmon plots to do this, and will be asking for help with it.
  • We must be able to confirm if there's an asymmetry caused by triggering in the 1Zk. For this to happen we need the trigger to be inefficient for Bs, but we should also provide a tool to emulate the effect of the 1Zk -> 1Z' which will show up any charge bias.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.